# Learning Prehensile Dexterity by Imitating and Emulating State-only Observations

### v1

@@ Abstract
When humans learn physical skills (e.g., learn to play tennis), we tend to first observe and learn what an expert is doing. But this is often insufficient. Therefore, we subsequently engage in practice, where we try to emulate the expert. Inspired by this observation, we introduce Combining IMitation and Emulation for Motion Refinement (CIMER) – a two-stage framework to learn dexterous prehensile manipulation skills from state-only observations. CIMER’s first stage involves imitation: simultaneously encode the complex interdependent motions of the robot hand and the object in a structured dynamical system. This results in a reactive motion generation policy that provides a reasonable motion prior, but lacks the ability to reason about contact effects due to the lack of action labels. The second stage involves emulation: learn a motion refinement policy to make adjustments to the motion prior of the robot hand such that the desired object motion is reenacted. CIMER is both task-agnostic (no task-specific reward design or shaping) and intervention-free (no need for additional teleoperated or labeled demonstrations). Detailed experiments reveal that i) Imitation alone is insufficient, but adding emulation drastically improves performance, ii) CIMER outperforms existing methods in terms of sample efficiency and the ability to generate realistic and stable motions, iii) CIMER can either zero-shot generalize or learn to adapt to novel objects from the YCB dataset, even outperforming expert policies trained with action labels in most cases.
@@ I Introduction
Learning dexterous manipulation skills involving multi-finger hands presents numerous challenges, such as high-dimensional state and action spaces, complex nonlinear dynamics, and contact effects. Unfortunately, when employed alone, both imitation learning (IL) and reinforcement learning (RL) tend to struggle with such complex high-dimensional tasks. Specifically, IL suffers from distribution shift [1, 2] and RL from poor sample efficiency [3]. To address these challenges, prior works have had success in leveraging demonstrations to expedite reinforcement learning [4, 5, 6, 7, 8, 9].
@@
However, collecting demonstrations for dexterous manipulation in the form of state-action pairs can be challenging and cumbersome due to complex system design and setup [10, 11]. To alleviate this burden, recent research has focused on learning dexterous manipulation skills from state-only observations [12, 13, 14]. However, learning purely from action-free data can result in kinematic motions that are oblivious to force and contact. This lack of knowledge is particularly limiting in dexterous prehensile manipulation tasks (e.g., grasping and tool use) as they exhibit heightened sensitivity to the applied force [12, 15, 16, 17]. To address this limitation, existing methods often require additional teleoperation demonstrations [14, 18], human-in-the-loop corrections [19], task-specific rewards [5, 13, 20, 21], or user-defined sub-goal images [7, 22].
@@
In this work, we contribute a novel task-agnostic framework to learn dexterous prehensile manipulation skills from state-only observations in two stages. We call this framework CIMER, short for Combining IMitation and Emulation for Motion Refinement (see Fig. 1). The first stage is Imitation: CIMER encodes the interdependent desired motions of both the robot hand and the object from the observations into a time-invariant dynamical system that acts as reactive motion generation policy. The second stage is Emulation: CIMER learns a motion refinement policy to adjust the generated robot motion by optimizing a task-agnostic reward that compensates for the lack of action information.
@@
Figure 1: CIMER is a task-agnostic and intervention-free framework to learn dexterous manipulation skills from state-only observations by learning to first generate interdependent desired motions of the robot hand and the object (Imitation), and then refine the generated robot motion in order to reenact the desired object motion (Emulation).
@@
Three key insights motivate our design of CIMER. First, dexterous prehensile manipulation involves complex and inherently interdependent motions of the robot hand and the object. As such, CIMER’s imitation stage captures both their motions simultaneously using a structured dynamical system (see Sec. III-C). Second, state-only observations of successful task executions contain valuable information about the robot hand’s motion, and importantly, unambiguously demonstrate how the object is supposed to move. The lack of action labels only impacts the motion of the robot hand, but not the object. As such, CIMER learns to refine the robot hand motion via reinforcement to ensure that the refined hand motion emulates the desired object motion (see Sec. III-D). Third, separating motion generation and refinement amounts to first learning what one must do (which can be done from state-only observations) and then learning how to do it (which can be learned by self-guided experimentation). Hence, unlike prior methods, CIMER neither requires in-context teleoperated demonstrations (intervention-free) nor task-specific reward designs (task-agnostic). We refer the reader to Sec. II for a detailed and nuanced discussion of related work.
@@
First, we demonstrate that directly tracking motions learned from state-only observations (i.e., without refinement) using a standard PD controller is rarely successful and necessitates meticulous expert tuning when it does. But CIMER achieves significant improvements in task performance, even while using the same PD controller across different tasks. We also uncover qualitative insights into how CIMER refines motions and why they succeed.
Second, we demonstrate that CIMER tends to achieve better sample efficiency compared to existing methods that learn purely from state-only observations. Moreover, we highlight CIMER’s ability to effectively leverage motion priors from observations to produce realistic behaviors, in contrast to baselines that tend to exploit the simulator and generate unrealistic behaviors.
Third, we show that CIMER significantly outperforms the baselines in terms of zero-shot generalization to unseen objects from the YCB dataset [23]. Notably, CIMER surpasses even the expert policy trained with action labels in many instances. We also demonstrate that CIMER is significantly faster in learning to work with new objects (i.e., skill transfer) for which it initially lacks good zero-shot generalization.
In summary, we contribute a task-agnostic and intervention-free framework called CIMER to learn dexterous prehensile manipulation skills from state-only observations. CIMER is task-agnostic as it can learn to effectively manipulate a variety of objects without relying on bespoke task-specific rewards. CIMER is intervention-free since it does not rely on in-situ expert demonstrations or teleoperation. Further, CIMER learns to ultimately produce PD targets instead of torques, which both promotes safety and facilitates hardware deployment as most multi-fingered hands only support position-based control.
@@ II Related Work
Learning dexterity from videos: Given the abundance of video data on the internet and the ease of data collection, recent research has focused on learning dexterous manipulation skills from videos. Indeed, videos are a natural source of state-only observations. A key challenge in learning from videos or state-only observations is the lack of data that speaks to the interaction of the robot and the object or the environment. Existing works circumvent this challenge in one of two ways. First, robots can be given an opportunity to learn from interactions, either in simulation [5, 9, 24, 25] or in the real world [7, 8, 22], in addition to videos. Second, robots can leverage additional in-domain teleoperated demonstrations from an expert [6, 14, 18, 20, 26]. Our work falls in the first category. Some of these prior works are limited to replicating a specific and constrained hand motion from a video clip, such as object grasping [8, 9, 24, 25], bagel flipping [6] or bottle opening [26]. While the other methods are capable of generating a variety of desired hand motions and can handle various object targets, they require one of the following: task-specific reward engineering [5, 20], user-provided sub-goal images as reward signals [7, 22], refinements tailored for dexterous grasping [12, 27], additional teleoperated robot demonstrations [14, 18]. In contrast, CIMER learns to generate a wide range of desired hand motions, without relying either on task-specific reward designs or expert intervention.
@@
Dexterity from teleoperation: A challenge related to learning from state-only observations is that of enabling a human operator to teleoperate a dexterous robotic hand [11, 28, 29, 30]. However, it is important to note that these methods are not learning autonomous policies. Instead, they leverage the real-time corrections and guidance provided by the human operator who monitors the robot’s behavior and compensates for its lack of knowledge about contact effects [31].
@@
Learning low-dim skills from videos: There is a large body of work that focuses on learning polices for low-dimensional manipulators (serial-link robots with parallel jaw grippers) from videos. One group of methods utilize a hierarchical framework to learn task-specific waypoint trajectories from human videos that can then be tracked using standard controllers [32, 33, 34, 35]. Another group learn a high-level planner that operates in a latent space and then use it to condition a low-level planner that can produce diverse robot behaviors [36, 37, 38]. However, unlike the low-dimensional skills that are handled by these methods, dexterous manipulation with multi-fingered hands inherently involves more complex finger-object interactions and contact effects. Further, similar to their counterparts that learn dexterous skills, many of these methods also rely on additional teleoperated demonstrations.
@@
Learning locomotion from observations: Our technical approach is inspired by research in learning locomotion skills. Prior work has demonstrated that robots can acquire locomotion skills through imitating real animals [39] or animated characters [40, 41]. These methods first extract the reference motion from video clips, and to account for foot-ground contact effects, refine the learned motions such that the resulting gaits closely track the demonstrated ones. However, given the object-centric nature of dexterous manipulation tasks, a reward strategy centered solely on imitating the observed robot motion may not necessarily ensure task success. This is precisely why CIMER involves an emulation stage where we refine the robot hand motion in order to track the desired object motions.
@@ III Approach
We begin by formulating the problem of learning dexterous manipulation skills from state-only observations, and then supply the details of our proposed solution.
@@ III-B Solution Overview
Our framework to address the above problem, Combining IMitation and Emulation for Motion Refinement (CIMER), operates in two stages: i) Imitation: learn a Motion Generation Policy $\Phi$ in the form of a structured dynamical system that encodes the desired reference motions for both the robot and the object from state-only observations, and ii) Emulation: learn a Motion Refinement policy $\Psi$ that compensates for the lack of action information by adjusting the generated reference motions such that the desired object motion is reenacted. The overall pseudo-code is given in Appendix  C.
@@ III-C Motion Generation via Imitation
We view the complex interdependent desired motions of the robot and the object as solutions to a learnable underlying nonlinear behavioral dynamical system [2, 42, 43]. In recent work, we developed a framework named KODex [44] that learns such underlying dynamics in a highly computationally-efficient manner using Koopman operator theory. Building on this, CIMER employs Koopman operator theory to learn the Motion Generation Policy $\Phi$ from the dataset $\mathcal{D}$. While KODex relies on demonstrations with action labels, CIMER does not. Formally, we formulate the $\Phi$ as a linear dynamical system in the lifted space: $g({\mathrm{h}}_{t+1},{\mathrm{o}}_{t+1})=\mathbf{K}\ \phi({\mathrm{h}}_{t},{\mathrm{o}}_{t})$, where $\phi:\mathbb{R}^{n+m}\rightarrow\mathbb{R}^{p}$ (with $p>(n+m)$) is a user-defined function that “lifts” the robot hand and object states to a higher-dimensional space in which the underlying dynamics appear linear [45], and $\mathbf{K}$ is the Koopman matrix. We use a task-agnostic second-order polynomial lifting function for all tasks in all our experiments. A key benefit of this approach is that we can use least squares to analytically compute $\mathbf{K}$ from $\mathcal{D}$ such that $\Phi$ effectively predicts the time evolution of the lifted states. To extract the original states without a decoder, we augment the original states to the lifted states before learning the Koopman operator. See Appendix D for details. We then use the learned dynamical system as an autonomous policy to generate reference hand motions for any initial condition.
@@ III-D Motion Refinement via Emulation
Merely tracking the hand trajectory generated by the motion generation policy $\Phi$ does not guarantee task success in dexterous prehensile manipulation tasks due to the small margin of error. Even slight deviations from the expert hand motion can have a drastic impact on the object (see Sec. IV-B for empirical validation). We address this issue by letting the robot “practice” the skill learned during imitation by emulating the desired object motion. Specifically, CIMER learns to refine finger and palm trajectories such that the robot reenacts the desired object motion. Interacting with the object helps compensate for the lack of action labels and account for contact effects. Once learned, CIMER employs a task-agnostic PD controller to track the refined hand motions.
@@
We formulate emulation as a reinforcement learning problem. We train a Motion Refinement Policy $\Psi$ to maximize a reward function that captures the essence of emulation: reenacting the desired object motion. To reduce the variance in the training, we use Generalized Advantage Estimator (GAE) [46] and PPO [47].
@@
Figure 2: CIMER refines the motion prior based on the context.
@@
State & Action Space: We parameterize $\Psi$ as an MLP network that predicts the current action: $\hat{\mathrm{h}}_{t}=\Psi(\mathrm{q}_{t},\mathrm{g}_{t})$. Here, $\hat{\mathrm{h}}(t)\in\mathcal{H}$ contains the refined PD targets of the robot hand at time $t$, $\mathrm{q}_{t}=(\bar{\mathrm{h}}_{t-2:t},\bar{\mathrm{o}}_{t-2:t},\hat{\mathrm{h}}_{t-3:t-1})$ is the context containing the history of hand and object states and refined PD targets, and $\mathrm{g}_{t}=(\mathrm{h}_{t+1,t+5,t+10},\mathrm{o}_{t+1,t+5,t+10})$ provides the motion prior (predicted by the motion generation policy $\Phi$) that needs to be refined. This design allows CIMER to reason about action effects when refining the hand motion generated by the motion generator. See Fig. 2 for an illustration.
@@
Rewards: Unlike prior works that rely on task-specific reward design and shaping, we use tracking errors of both the robot hand and the object as reward signals. Formally, we define the reward function as $r_{t}=r_{t}^{h}+r_{t}^{o}+r_{t}^{b}$, where ${r_{t}^{h}=\exp\left(-k^{h}||\bar{\mathrm{h}}_{t+1}-\mathrm{h}_{t+1}||^{2}\right)}$, ${r_{t}^{o}=\exp\left(-k^{o}||\bar{\mathrm{o}}_{t+1}-\mathrm{o}_{t+1}||^{2}\right)}$, $r_{t}^{b}$ is a bonus reward that is trigger when the object tracking error is smaller than a threshold $\epsilon_{o}$. Here, $\bar{\mathrm{h}}_{t+1},\mathrm{h}_{t+1}$, and $\bar{\mathrm{o}}_{t+1},\mathrm{o}_{t+1}$ denote the hand and object states from both the robot and reference motion, respectively. Note that $r_{t}^{h}$ minimizes deviation from the reference hand motion $\mathrm{h}_{t+1}$ while $r_{t}^{o}$ encourages the robot hand to emulate the reference object motion $\mathrm{o}_{t+1}$.
@@
Warm Start: Since the motion generation policy $\Phi$ provides a reasonable motion prior, we expect the motion refinement policy $\Psi$ to only make local adjustments and account for contact effects. As such, we use the reference motions from $\Phi$ to warm start $\Psi$ by minimizing a behavior cloning loss. To counter covariate shift, we inject random noise into the policy’s input and use the same outputs as corrective labels [48, 49, 50]. This warm start represents yet another way in which the motion prior influences the emulation process (in addition to via the reward function and the policy input).

### v3

@@ Abstract
When human acquire physical skills (e.g., tool use) from experts, we tend to first learn from merely observing the expert. But this is often insufficient. We then engage in practice, where we try to emulate the expert and ensure that our actions produce similar effects on our environment. Inspired by this observation, we introduce Combining IMitation and Emulation for Motion Refinement (CIMER) – a two-stage framework to learn dexterous prehensile manipulation skills from state-only observations. CIMER’s first stage involves imitation: simultaneously encode the complex interdependent motions of the robot hand and the object in a structured dynamical system. This results in a reactive motion generation policy that provides a reasonable motion prior, but lacks the ability to reason about contact effects due to the lack of action labels. The second stage involves emulation: learn a motion refinement policy via reinforcement that adjusts the robot hand’s motion prior such that the learned object motion is reenacted. CIMER is both task-agnostic (no task-specific reward design or shaping) and intervention-free (no additional teleoperated or labeled demonstrations). Detailed experiments with prehensile dexterity reveal that i) imitation alone is insufficient, but adding emulation drastically improves performance, ii) CIMER outperforms existing methods in terms of sample efficiency and the ability to generate realistic and stable motions, iii) CIMER can either zero-shot generalize or learn to adapt to novel objects from the YCB dataset, even outperforming expert policies trained with action labels in most cases. Source code and videos are available at https://sites.google.com/view/cimer-2024/.
@@ I Introduction
Learning dexterous manipulation skills involving multi-finger hands (e.g., relocation and tool use) presents numerous challenges, such as high-dimensional state and action spaces, complex nonlinear dynamics, and contact effects. Unfortunately, both imitation learning (IL) and reinforcement learning (RL) tend to struggle with such complex tasks when employed alone. Specifically, IL suffers from distribution shift [1, 2] and RL from poor sample efficiency [3]. To address these challenges, prior works have shown promise in leveraging demonstrations to expedite reinforcement learning [4, 5, 6, 7, 8, 9].
@@
However, collecting demonstrations for dexterous manipulation in the form of state-action pairs can be challenging and cumbersome due to complex system design and setup [10, 11]. To alleviate this burden, recent research has focused on learning dexterous manipulation skills from state-only observations [12, 13, 14]. However, learning without action labels can result in kinematic motions that are oblivious to force and contact. This lack of knowledge is particularly limiting in dexterous prehensile manipulation tasks (e.g., grasping and tool use) as they exhibit heightened sensitivity to applied force [12, 15, 16, 17]. To compensate, existing methods often require additional teleoperated demonstrations [14, 18], human-in-the-loop corrections [19], task-specific rewards [5, 13, 20, 21], or user-defined sub-goal images [7, 22].
@@
In this work, we contribute a novel two-stage framework to learn dexterous prehensile manipulation skills from state-only observations. We call our framework CIMER, short for Combining IMitation and Emulation for Motion Refinement (see Fig. 1). The first stage is Imitation: CIMER encodes the interdependent desired motions of both the robot hand and the object from the observations into a structured time-invariant dynamical system that acts as a reactive policy and a reasonable motion prior (see Sec. III-C). The second stage is Emulation: CIMER learns to refine the robot’s motion prior based on context by optimizing a task-agnostic reward that incentivizes the reenactment of the learned object motion [23].
@@
Figure 1: CIMER is a task-agnostic and intervention-free framework to learn dexterous manipulation skills from state-only observations by learning to first generate interdependent desired motions of the robot hand and the object (Imitation), and then refine the generated robot motion in order to reenact the learned object motion (Emulation).
@@
Three key insights motivate our design of CIMER. First, dexterous prehensile manipulation involves complex and inherently interdependent motions of the robot hand and the object. As such, CIMER’s imitation stage captures both their motions simultaneously using a single but structured dynamical system. Second, state-only observations of the expert unambiguously demonstrate how the object is supposed to move, and contain valuable (albeit crude) information about the robot hand’s motion. As such, CIMER learns to only refines the hand motion and relies on the learned object motion for reward signals. Third, separating motion generation and refinement amounts to first learning what the robot must do from state-only observations, and then learning how to do it via self-guided experimentation. Unlike prior methods, CIMER is both intervention-free and task-agnostic – it neither requires in-context additional demonstrations nor task-specific reward design. See Sec. II for a detailed discussion of related work.
@@
First, we demonstrate that pure imitation (without emulation) is rarely successful, and necessitates meticulous expert tuning of the low-level controller when it does. But combining imitation and emulation (i.e., CIMER) significantly improves the task success rate, even while using the same controller across tasks. We also uncover qualitative insights into how CIMER refines motions and why they succeed.
Second, we demonstrate that CIMER tends to achieve better sample efficiency compared to existing methods that can learn from state-only observations. Moreover, we highlight CIMER’s ability to effectively leverage motion priors (encoded in a dynamical system) to produce realistic and stable motions, in contrast to baselines that tend to exploit the simulator and generate aggressive behaviors. We also show that policies trained using CIMER exhibit better robustness against changes to physical parameters such as mass and damping.
Third, we show that CIMER significantly outperforms the baselines in terms of zero-shot generalization to 17 unseen objects from the YCB dataset [24]. Notably, CIMER surpasses even the expert policy trained with action labels in many instances. When CIMER can’t readily generalize to a novel object with significant differences, it can effectively leverage previously learned skills to expedite adaptation.
In summary, we contribute CIMER – a two-stage framework that effectively combines imitation and emulation in order to learn dexterous prehensile manipulation skills from state-only observations. CIMER first learns a reasonable motion prior by imitating observations without relying on action labels or user interventions, and then employs a task-agnostic and self-guided strategy to learn how to refine the hand’s motion so that it emulates the learned object motion.
@@ II Related Work
Learning dexterity from observations: Given the abundance of video data on the internet and the advances in motion retargetting (e.g., [5]), recent research has focused on learning dexterous manipulation skills from videos or state-only observations in general. A key challenge in learning from videos or state-only observations is the lack of action labels that speaks to the contact effects between the robot, the object, and the environment. Existing works circumvent this challenge in one of two ways. Some methods provide robots an opportunity to learn from interactions, either in simulation [5, 9, 25, 26] or in the real world [7, 8, 22]. Other methods leverage additional in-domain teleoperated demonstrations from an expert [6, 14, 18, 20, 27]. Our work falls under the first category since it does not rely on any additional demonstrations. Unlike CIMER, many of the prior methods in the first category are limited to replicating a specific hand motion from a video clip, such as object grasping [8, 9, 25, 26], bagel flipping [6] or bottle opening [27]. While the remaining methods are capable of learn a variety of skills and can handle changes to the objects’ initial and target configurations, they require at least one of the following: task-specific reward engineering [5, 20], user-provided sub-goal images as reward signals [7, 22], and bespoke refinements tailored to dexterous grasping [12, 28]. In contrast, CIMER can learn a variety of skills requiring neither task-specific reward shaping nor expert interventions.
@@
Learning dexterous teleoperation: Another class of existing methods enable a human operator to seamlessly teleoperate a dexterous robotic hand by leveraging learning to close the embodiment gap [11, 29, 30, 31]. However, it is important to note that these methods do not learn autonomous policies. Instead, they rely on real-time corrections and guidance from the human operator who tends to compensate for the robot’s inability to reason about contact effects [32].
@@
Learning low-dim skills from videos: There is a large body of work that focuses on learning polices for low-dimensional manipulators (serial-link robots with parallel jaw grippers) from videos. One group of methods utilize a hierarchical framework to learn task-specific waypoint trajectories from human videos that can then be tracked using standard controllers [33, 34, 35, 36]. Other methods learn a high-level planner that operates in a latent space and then use it to condition a low-level planner that can produce diverse robot behaviors [37, 38, 39]. However, unlike the low-dimensional skills learned by these methods, dexterous manipulation with multi-fingered hands inherently involves more complex finger-object interactions and contact effects. Further, similar to their counterparts that learn dexterous skills, many of these methods also rely on additional teleoperated demonstrations.
@@
Learning locomotion from observations: Our approach to emulation is inspired by research in learning locomotion skills by imitating animals [40] or animated characters [41, 42]. These methods first extract the reference motion from video clips, and refine the learned motions to account for foot-ground contact effects and to ensure that the resulting gaits closely track the demonstrated ones. However, merely imitating the observed robot motion is unlikely to succeed in dexterous manipulation since it is object-centric (see Sec. IV-B for empirical validation). In contrast, we emphasize the reenactment of the learned object motion.
@@ III Approach
We begin by formulating the problem of learning dexterous manipulation skills from state-only observations.
@@ III-B Solution Overview
Our framework to address the above problem, Combining IMitation and Emulation for Motion Refinement (CIMER), operates in two stages: i) Imitation: learn a Motion Generation Policy $\Phi$ in the form of a structured dynamical system that encodes the desired reference motions for both the robot and the object from state-only observations, and ii) Emulation: learn a Motion Refinement policy $\Psi$ that compensates for the lack of action information by refining the desired robot motions such that the learned object motion is achieved. The overall pseudo-code is given in Appendix  D.
@@ III-C Motion Generation via Imitation
We view the complex interdependent desired motions of the robot and the object as solutions to a learnable underlying nonlinear behavioral dynamical system [2, 43, 44]. Once learned, these dynamical systems can be integrated to predict the next desired robot state from initial conditions or the current state [2]. In recent work, we developed a framework named KODex [45] that learns such underlying dynamics in a highly computationally-efficient manner using Koopman operator theory. Data-driven Koopman-based approaches help effectively encode highly nonlinear dynamical systems in a linear system using lifted global linearization [46, 47]. Building on this, CIMER employs a Koopman-based method theory to learn the Motion Generation Policy $\Phi$ from the dataset $\mathcal{D}$. Note that KODex requires access to action labels, but CIMER does not. Formally, we formulate the $\Phi$ as a linear dynamical system in the lifted space: $\phi({\mathrm{h}}_{t+1},{\mathrm{o}}_{t+1})=\mathbf{K}\ \phi({\mathrm{h}}_{t},{\mathrm{o}}_{t})$, where $\phi:\mathbb{R}^{n+m}\rightarrow\mathbb{R}^{p}$ (with $p>(n+m)$) is a user-defined function that “lifts” the robot hand and object states to a higher-dimensional space in which the underlying dynamics appear linear, and $\mathbf{K}$ is the Koopman matrix [46, 47]. We use a task-agnostic second-order polynomial lifting function for all tasks in all our experiments. A key benefit of using a Koopman-based approach is that we can analytically compute $\mathbf{K}$ (and thereby our motion generation policy $\Phi$) from $\mathcal{D}$ using least squares [47]. To extract the original states without a decoder, we augment the original states to the lifted states before learning the Koopman operator. See Appendix E for details. We then use the learned dynamical system as an autonomous policy to predict desired hand and object motions for any initial condition.
@@ III-D Motion Refinement via Emulation
Merely tracking the hand trajectory generated by the motion generation policy $\Phi$ does not guarantee task success in dexterous prehensile manipulation tasks due to the small margin for error and the lack of reasoning about contact effects (see Sec. IV-B for empirical validation). As such, CIMER refines the robot’s hand motion so as to ensure emulation of the learned object motion. Specifically, CIMER learns $\Psi$ to refine finger and palm trajectories (as opposed to all the DoFs) generated by $\Phi$ such that the robot reenacts the object motion generated by $\Phi$. Interacting with the object helps compensate for the lack of action labels and account for contact effects. Once learned, CIMER employs a task-agnostic PD controller to track the refined hand motions.
@@
We formulate emulation as an RL problem and maximize rewards computed from policies learned via imitation. Our task-agnostic reward function captures the essence of emulation: reenacting the desired object motion rather than only the expert’s motion [23]. To reduce training variance, we use Generalized Advantage Estimator (GAE) [48] and PPO [49].
@@
Figure 2: CIMER generates a motion prior based on initial conditions and refines it based on context to generate PD targets for the hand.
@@
State & Action Space: We parameterize $\Psi$ as an MLP network that predicts the current action: $\hat{\mathrm{h}}_{t}=\Psi(\mathrm{q}_{t},\mathrm{g}_{t})$. Here, $\hat{\mathrm{h}}(t)\in\mathcal{H}$ contains the refined PD targets of the robot hand at time $t$, $\mathrm{q}_{t}=(\bar{\mathrm{h}}_{t-2:t},\bar{\mathrm{o}}_{t-2:t},\hat{\mathrm{h}}_{t-3:t-1})$ is the context containing the history of hand and object states and refined PD targets, and $\mathrm{g}_{t}=(\mathrm{h}_{t+1,t+5,t+10},\mathrm{o}_{t+1,t+5,t+10})$ provides the motion prior (predicted by the motion generation policy $\Phi$) that needs to be refined. Since the necessary refinements are likely to be local and fine-grained, we restrict $\Psi$ to only refine the finger and palm trajectories. See Fig. 2 for an illustration.
@@
Rewards: Unlike prior works that rely on task-specific reward shaping, we use tracking errors of both the robot hand and the object as reward signals. Formally, we define the reward function as $r_{t}=r_{t}^{h}+r_{t}^{o}+r_{t}^{b}$, where ${r_{t}^{h}=\exp\left(-k^{h}||\bar{\mathrm{h}}_{t+1}-\mathrm{h}_{t+1}||^{2}\right)}$, ${r_{t}^{o}=\exp\left(-k^{o}||\bar{\mathrm{o}}_{t+1}-\mathrm{o}_{t+1}||^{2}\right)}$, $r_{t}^{b}$ is a bonus reward that is triggered when the object tracking error is smaller than a threshold $\epsilon_{o}$. Here, $\bar{\mathrm{h}}_{t+1},\mathrm{h}_{t+1}$, and $\bar{\mathrm{o}}_{t+1},\mathrm{o}_{t+1}$ denote the hand and object states from both the robot and reference motion, respectively. Note that $r_{t}^{h}$ minimizes deviation from the reference hand motion $\mathrm{h}_{t+1}$ while $r_{t}^{o}$ encourages the robot hand to emulate the reference object motion $\mathrm{o}_{t+1}$. We include the training parameters in Appendix B.
@@
Warm Start: Since the motion generation policy $\Phi$ provides a reasonable motion prior, we expect the motion refinement policy $\Psi$ to only make local adjustments and account for contact effects. As such, we use the reference motions from $\Phi$ to warm start $\Psi$ by minimizing a behavior cloning loss. To counter covariate shift, we inject random noise into the policy’s input and use the same outputs as corrective labels [50, 51, 52]. This warm start represents yet another way in which imitation influences the emulation process (in addition to via the reward function and the policy input).